# GeoDynamics: A Geometric State‑Space Neural Network for Understanding Brain Dynamics on Riemannian Manifolds

### v1

@@ 4.3 Ablation Studies
Multi‐class classification (ADNI dataset): We extended the binary classification task to multiple clinical stages. As reported in Table 2, GeoDynamics consistently achieves higher accuracy, precision, and F1-score across all classes compared with both GNN- and sequence-based baselines. This demonstrates robust generalization to more challenging multi-class settings.
@@ A.1 Literature Survey
Recurrent neural networks (RNNs) have been reformulated as ordinary differential equations (ODEs) with continuous-time system states, as highlighted by LTCNet [30]. These models serve as effective algorithms for modeling time series data and are widely utilized across medical, industrial, and business domains. For instance, [9] has demonstrated its potential for brain state recognition and [29] achieves continuous modeling of dynamic brain signals using ODEs. Furthermore, the survey proposed by [55] provides a comprehensive overview of ODE applications in the field of medical imaging, showcasing their practicality and impact in this domain. Following this, several manifold-based RNN models have emerged. For instance, [10] introduced a statistical recurrent model defined on the manifold of symmetric positive definite (SPD) matrices and evaluated its diagnostic potential for neuroimaging applications. This approach underscores the effectiveness of utilizing manifold-based techniques to enhance the performance of RNNs in complex medical contexts. The RNN model formulated on Riemannian manifolds [37] is robustly supported by mathematical theory, as it utilizes covariance information to dynamically model time-series data [38]. This capability allows it to capture richer and more subtle representations within a higher-dimensional latent space. Such an approach is particularly effective in modeling complex data structures, such as capturing the functional dynamics [15; 33], where the relationships among data points are inherently geometric. By operating within the manifold framework, these models adeptly accommodate the intricacies of underlying data distributions, thereby enhancing both interpretability and predictive performance.
@@ A.5 SPD Padding $\cdot$] Operation
Given a SPD matrices $X\in Sym_{N}^{+}$ and a small positive value $\theta$, the assemble matrix $Y=diag(\theta,X)=\left[\begin{array}[]{cc}\theta&0\\ 0&X\end{array}\right]$ is a SPD matrix.
Proof: First, $Y$ is a symmetric, since $Y^{\top}=\begin{bmatrix}\theta&0\\ 0&X\end{bmatrix}^{\top}=\begin{bmatrix}\theta&0\\ 0&X\end{bmatrix}=Y$. Then, to show that $Y$ is positive definite, we need to verify that for any non-zero vector $z=\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}\in\mathbb{R}^{N+1}$, the quadratic form $z^{\top}Yz$ is strictly positive.
@@
$z^{\top}Yz=\begin{bmatrix}z_{1}&z_{2}^{\top}\end{bmatrix}\begin{bmatrix}\theta&0\\ 0&X\end{bmatrix}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=z_{1}^{2}\theta+z_{2}^{\top}Xz_{2}.$ (36)
Since $\theta>0$, the term $z_{1}^{2}\theta\geq 0$, and it is strictly positive if $z_{1}\neq 0$.
@@ A.10 Limitations and Future Work
Second, the effectiveness of manifold-based modeling depends on the assumption that input SPD matrices lie within a well-behaved region of the manifold (e.g., within a geodesic ball), which ensures the uniqueness and stability of the Fréchet mean. When this assumption is violated—such as in highly noisy or poorly conditioned data (rank-deficient), the performance of the method may degrade.

### v2

@@ 4.3 Ablation Studies
Multi‐class classification (ADNI dataset): We extended the binary classification task to multiple clinical stages. As reported in Table 2, GeoDynamics consistently achieves higher accuracy, precision, and F1-score across all classes compared with both spatial- and sequence-based baselines. This demonstrates robust generalization to more challenging multi-class settings.
@@ A.1 Literature Survey
Recurrent neural networks (RNNs) have been reformulated as ordinary differential equations (ODEs) with continuous-time system states, as highlighted by LTCNet [30]. These models serve as effective algorithms for modeling time series data and are widely utilized across medical, industrial, and business domains. For instance, [9] has demonstrated its potential for brain state recognition and [29] achieves continuous modeling of dynamic brain signals using ODEs. Furthermore, the survey proposed by [54] provides a comprehensive overview of ODE applications in the field of medical imaging, showcasing their practicality and impact in this domain. Following this, several manifold-based RNN models have emerged. For instance, [10] introduced a statistical recurrent model defined on the manifold of symmetric positive definite (SPD) matrices and evaluated its diagnostic potential for neuroimaging applications. This approach underscores the effectiveness of utilizing manifold-based techniques to enhance the performance of RNNs in complex medical contexts. The RNN model formulated on Riemannian manifolds is robustly supported by mathematical theory, as it utilizes covariance information to dynamically model time-series data [37]. This capability allows it to capture richer and more subtle representations within a higher-dimensional latent space. Such an approach is particularly effective in modeling complex data structures, such as capturing the functional dynamics [15; 33], where the relationships among data points are inherently geometric. By operating within the manifold framework, these models adeptly accommodate the intricacies of underlying data distributions, thereby enhancing both interpretability and predictive performance.
@@ A.5 SPD Padding $\cdot$] Operation
Given a SPD matrices $X\in Sym_{N}^{+}$ and a small positive value $\rho$, the assemble matrix $Y=diag(\rho,X)=\left[\begin{array}[]{cc}\rho&0\\ 0&X\end{array}\right]$ is a SPD matrix.
Proof: First, $Y$ is a symmetric, since $Y^{\top}=\begin{bmatrix}\rho&0\\ 0&X\end{bmatrix}^{\top}=\begin{bmatrix}\rho&0\\ 0&X\end{bmatrix}=Y$. Then, to show that $Y$ is positive definite, we need to verify that for any non-zero vector $z=\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}\in\mathbb{R}^{N+1}$, the quadratic form $z^{\top}Yz$ is strictly positive.
@@
$z^{\top}Yz=\begin{bmatrix}z_{1}&z_{2}^{\top}\end{bmatrix}\begin{bmatrix}\rho&0\\ 0&X\end{bmatrix}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix}=z_{1}^{2}\rho+z_{2}^{\top}Xz_{2}.$ (36)
Since $\rho>0$, the term $z_{1}^{2}\rho\geq 0$, and it is strictly positive if $z_{1}\neq 0$.
@@ A.10 Limitations and Future Work
Second, the effectiveness of manifold-based modeling depends on the assumption that input SPD matrices lie within a well-behaved region of the manifold (e.g., within a geodesic ball), which ensures the uniqueness and stability of the Fréchet mean. When this assumption is violated, such as in highly noisy or poorly conditioned data (rank-deficient), the performance of the method may degrade.